An aqueous solution of perchloric acid is standardized by titration with a 0.193 M solution of calcium hydroxide.

If 20.6 mL of base are required to neutralize 27.0 mL of the acid, what is the molarity of the perchloric acid solution?

__________ M perchloric acid
1
Expert's answer
2019-03-18T16:22:54-0400

"c(HClO4)=\dfrac{c(Ca(OH)2)*V(Ca(OH)2)}{V(HClO4)}=\dfrac{0.193M*20.6mL}{27mL}=0.147 mol/L"